# OptiMindTune: A Multi-Agent Framework for Intelligent Hyperparameter Optimization

### v1

@@ Abstract
Hyperparameter optimization (HPO) is a critical yet challenging aspect of machine learning model development, significantly impacting model performance and generalization. Traditional HPO methods often struggle with high dimensionality, complex interdependencies, and computational expense. This paper introduces OptiMindTune, a novel multi-agent framework designed to intelligently and efficiently optimize hyperparameters. OptiMindTune leverages the collaborative intelligence of three specialized AI agents – a Recommender Agent, an Evaluator Agent, and a Decision Agent – each powered by Google’s Gemini models. These agents address distinct facets of the HPO problem, from model selection and hyperparameter suggestion to robust evaluation and strategic decision-making. By fostering dynamic interactions and knowledge sharing, OptiMindTune aims to converge to optimal hyperparameter configurations more rapidly and robustly than existing single-agent or monolithic approaches. Our framework integrates principles from advanced large language models, and adaptive search to achieve scalable and intelligent AutoML. We posit that this multi-agent paradigm offers a promising avenue for tackling the increasing complexity of modern machine learning model tuning.
@@ I Introduction
This paper introduces OptiMindTune, a novel multi-agent framework for intelligent hyperparameter optimization. Unlike conventional HPO approaches that often operate as monolithic optimizers, OptiMindTune decentralizes the HPO process by employing a community of specialized, intelligent agents, each powered by Google’s Gemini models. Each agent is designed to contribute to the overall optimization goal by focusing on specific tasks: the Recommender Agent for model and hyperparameter suggestion, the Evaluator Agent for robust training and validation, and the Decision Agent for strategic search guidance. This collaborative and distributed intelligence aims to overcome the limitations of single-agent systems, enabling more efficient exploration, faster convergence, and greater robustness through an adaptive optimization loop.
@@
  • We propose OptiMindTune, a novel multi-agent framework for intelligent hyperparameter optimization, departing from traditional monolithic approaches and powered by advanced large language models.
@@ II-B4 Multi-Agent Systems for AutoML
The field of multi-agent AutoML is emerging, with initial efforts focusing on distributed approaches, such as systems utilizing multiple AutoML processes running in parallel, managed by a higher-level orchestrator for resource allocation or result aggregation [19]. However, these frameworks often lack the dynamic, intelligent collaboration required for sophisticated hyperparameter optimization (HPO). Our proposed OptiMindTune framework introduces a novel three-agent architecture—Recommender, Evaluator, and Decision—powered by Google’s Gemini models, enabling a dynamic optimization loop that achieves high accuracies. Its Decision agent explicitly balances exploration of diverse configurations and exploitation of promising ones, optimizing the search process, while real-time logging of all agent interactions provides a transparent audit trail for analysis and debugging. In contrast, AgentHPO [23] employs a two-agent system—Creator and Executor—powered by OpenAI’s GPT models, with iterative refinement based on experimental logs. OptiMindTune’s focused design and adaptive strategy highlight its potential for precise, resource-efficient HPO in classification tasks, complementing the broader applicability of AgentHPO and advancing the paradigm of intelligent, collaborative multi-agent AutoML.
@@ III Methodology
OptiMindTune is conceptualized as a distributed, intelligent multi-agent framework for hyperparameter optimization. The core principle is to decompose the complex HPO problem into manageable sub-tasks, each handled by a specialized agent powered by Google’s Gemini models, fostering a collaborative ecosystem that collectively drives towards optimal hyperparameter configurations for scikit-learn classifiers.
@@ III-A OptiMindTune Architecture
The OptiMindTune framework comprises three distinct types of AI agents, each with a specific role and responsibilities, interacting through a shared communication mechanism. (A system diagram illustrating the architecture will be provided in the final paper submission.)
  • Recommender Agent: This agent acts as the intelligent explorer and knowledge synthesizer. Powered by Google’s Gemini models, it analyzes dataset characteristics and reviews past performance history to suggest suitable machine learning models (e.g., from scikit-learn) and their corresponding hyperparameter configurations.
@@
    • Powered by: Google’s Gemini models for intelligent reasoning, pattern recognition, and suggestion generation based on contextual information.
@@
        Powered by: Google’s Gemini models for strategic reasoning, goal-oriented decision-making, and assessing trade-offs.
@@ III-B1 Communication Flow (Optimization Loop)
    • Recommender Suggests: The Recommender Agent, drawing upon its Gemini-powered intelligence, analyzes the dataset characteristics and the current optimization history. It then suggests a set of model and hyperparameter configurations, along with its reasoning, to the Decision Agent.
@@
    • Decision Agent Guides: The Decision Agent evaluates the reported performance. Using its Gemini-powered reasoning and the configured goals, it decides whether to accept the configuration (if it meets criteria), or to guide the Recommender Agent towards further exploration or exploitation. This includes determining if the search should terminate.
@@ IV Experiments
To rigorously evaluate the performance of OptiMindTune, we propose a comprehensive experimental setup comparing its efficacy against established HPO baselines across a variety of machine learning tasks and datasets, specifically for scikit-learn classifiers.
@@ IV-A Datasets
We will utilize a diverse set of public benchmark datasets to demonstrate OptiMindTune’s generalizability and robustness on classification tasks.
  • UCI Machine Learning Repository: We will select several well-known classification datasets from the UCI repository, such as Iris, Wine, Breast Cancer Wisconsin (Diagnostic), and Adult. These datasets represent different data types and complexities, allowing for a thorough evaluation of OptiMindTune on traditional classification tasks.
  • OpenML Datasets: OpenML [20] provides a vast collection of datasets and tasks, facilitating reproducible and large-scale experimentation. We will leverage OpenML classification datasets with varying numbers of features, instances, and classes to assess scalability. Specific datasets will be chosen to represent common classification challenges.
  • Tabular Benchmarks: For evaluating the hyperparameter optimization on more challenging tabular data, we will also consider tabular benchmarks designed for AutoML, such as those used in the OpenML-CC18 benchmark suite [21], specifically focusing on classification tasks.
@@ IV-B Machine Learning Models
OptiMindTune is designed to optimize hyperparameters for scikit-learn classifiers. We will apply OptiMindTune to optimize hyperparameters for a range of commonly used scikit-learn classification models:
  • Support Vector Classifiers (SVC)
  • Random Forest Classifiers
  • Logistic Regression
The Recommender Agent will suggest combinations of these models and their specific hyperparameters.
@@ IV-C Baseline Methods
We will compare OptiMindTune against several state-of-the-art and widely used HPO techniques, implemented via scikit-learn compatible wrappers or direct calls to their libraries:
  • Random Search [1]: A strong baseline for high-dimensional search spaces.
  • Bayesian Optimization (e.g., using Scikit-Optimize) [2]: A representative sequential model-based optimization method.
  • Optuna (TPE) [10]: A popular and efficient HPO framework based on the Tree-structured Parzen Estimator.
@@ IV-D Evaluation Metrics
The performance of HPO methods will be primarily evaluated based on:
  • Achieved Objective Value: The best performance (e.g., Accuracy) of the machine learning model found by the HPO algorithm within a given budget, specifically using cross-validation scores.
  • Convergence Speed: How quickly the HPO method finds near-optimal configurations. This will be measured by plotting the objective value against the number of evaluations or wall-clock time.
  • Computational Budget Efficiency: The total computational resources (e.g., CPU hours, or number of model evaluations including cross-validation folds) consumed to reach a certain performance level or throughout the entire HPO process.
  • Robustness: The consistency of performance across multiple independent runs and different datasets, indicating the reliability of the HPO method.
@@ IV-E Experimental Setup
  • Hardware: Experiments will be conducted on a cluster of machines with consistent CPU/GPU specifications to ensure fair comparison of wall-clock time.
  • Budget Allocation: Each HPO method will be allocated a fixed budget, either in terms of total evaluation time or maximum number of model evaluations (inclusive of cross-validation folds). This ensures a fair comparison of efficiency.
  • Cross-Validation: For each hyperparameter configuration, model performance will be evaluated using k-fold cross-validation (e.g., $k=5$ or $k=10$) within the Evaluator Agent, to obtain robust estimates of performance and minimize variance.
  • Reproducibility: All experiments will be run with fixed random seeds for reproducibility. We will perform multiple independent runs for each HPO method on each dataset to account for stochasticity inherent in some methods and agent interactions.
  • Hyperparameter Search Space Definition: For each model and dataset, a reasonable hyperparameter search space will be defined, consistent across all HPO methods. The Recommender Agent will operate within this defined space.
@@ V-D Discussion of Results
VI Experiments
VI-A Datasets
We used three standard UCI classification datasets:
  • Breast Cancer Wisconsin (Diagnostic): Binary classification.
  • Iris: Multi-class classification with three classes.
  • Wine: Multi-class classification with three classes.
VI-B Machine Learning Models
  • LogisticRegression: A linear model.
  • RandomForestClassifier: An ensemble tree-based model.
VI-C Baseline Methods
We compared OptiMindTune against:
VI-D Evaluation Metrics
Performance was assessed using:
  • Achieved Objective Value: Mean cross-validated accuracy.
VI-E Experimental Setup
  • Hardware: Consistent hardware was used.
  • Reproducibility: Fixed random seeds ensured consistency.
VII Discussion and Future Work
VII-A Reflection on Findings
The experimental results are expected to validate OptiMindTune’s efficacy as a multi-agent framework for intelligent hyperparameter optimization. We anticipate observing that OptiMindTune achieves competitive or superior performance compared to state-of-the-art baselines in terms of objective value, convergence speed, and computational efficiency across a diverse set of datasets and machine learning models. The strength of OptiMindTune lies in its ability to leverage the collective intelligence of specialized agents, each powered by Google’s Gemini models, allowing for dynamic adaptation to the HPO landscape. The clear division of labor among the Recommender, Evaluator, and Decision Agents, with their sophisticated communication and internal Gemini-driven reasoning, facilitates a more comprehensive and adaptive search strategy than monolithic approaches. The explicit real-time logging and conversation tracking also provide unprecedented transparency into the AutoML process.
VII-B Limitations
@@
  • Interpretability of LLM Decisions: While the agents provide reasoning, the internal workings of the Gemini models themselves remain a black box. Understanding precisely why a Gemini-powered agent makes a certain recommendation or decision can be challenging, impacting the overall interpretability of the HPO process.
@@
VII-C Potential Extensions and Future Work
@@
VIII Conclusion
Hyperparameter optimization remains a formidable challenge in machine learning, demanding intelligent and efficient solutions. This paper introduced OptiMindTune, a novel multi-agent framework that reimagines HPO as a collaborative effort among specialized intelligent agents, each powered by Google’s Gemini models. By decoupling responsibilities into distinct roles for recommendation, evaluation, and strategic decision-making, and fostering dynamic interactions, OptiMindTune aims to overcome the limitations of traditional monolithic approaches. The proposed architecture promises enhanced adaptability, robustness, and efficiency in navigating complex hyperparameter landscapes for scikit-learn classifiers. While initial conceptualization and experimental design lay a strong foundation, the future work outlined points towards exciting avenues for extending OptiMindTune to tackle even more challenging optimization problems in the rapidly evolving field of machine learning, especially in the context of deep learning and neural architecture search, leveraging the ever-increasing capabilities of advanced AI models. OptiMindTune represents a significant step towards truly intelligent and autonomous hyperparameter optimization.

### v2

@@ Abstract
Hyperparameter optimization (HPO) plays a vital role in enhancing machine learning model performance but remains challenging due to high-dimensional search spaces, complex interdependencies among parameters, and significant computational cost. In this paper, we present OptiMindTune, a novel multi-agent framework for HPO that employs three specialized agents: a Recommender Agent for generating candidate hyperparameters, an Evaluator Agent for assessing model performance, and a Decision Agent for orchestrating the search process. Each agent operates using large language models (LLMs) to facilitate reasoning, interaction, and adaptation across the optimization process. By enabling structured communication and role-specific behaviors, OptiMindTune aims to improve convergence efficiency and robustness over traditional single-agent or black-box optimization methods. Our framework integrates principles from advanced large language models, and adaptive search to achieve scalable and intelligent AutoML. We posit that this multi-agent paradigm offers a promising avenue for tackling the increasing complexity of modern machine learning model tuning.
@@ I Introduction
This paper introduces OptiMindTune, a novel multi-agent framework for intelligent hyperparameter optimization. Unlike conventional HPO approaches that often operate as monolithic optimizers, OptiMindTune decentralizes the HPO process by employing a community of specialized, intelligent agents, each powered by large language models. Each agent is designed to contribute to the overall optimization goal by focusing on specific tasks: the Recommender Agent for model and hyperparameter suggestion, the Evaluator Agent for robust training and validation, and the Decision Agent for strategic search guidance. This collaborative and distributed intelligence aims to overcome the limitations of single-agent systems, enabling more efficient exploration, faster convergence, and greater robustness through an adaptive optimization loop.
@@
  • We propose OptiMindTune, a novel multi-agent framework for intelligent hyperparameter optimization, departing from traditional monolithic approaches and powered by large language models.
@@ II-B4 Multi-Agent Systems for AutoML
The field of multi-agent AutoML is emerging, with initial efforts focusing on distributed approaches, such as systems utilizing multiple AutoML processes running in parallel, managed by a higher-level orchestrator for resource allocation or result aggregation [19]. However, these frameworks often lack the dynamic, intelligent collaboration required for sophisticated hyperparameter optimization (HPO). AgentHPO [23] employs a two-agent system—Creator and Executor—powered by OpenAI’s GPT models, with iterative refinement based on experimental logs. OptiMindTune’s focused design and adaptive strategy highlight its potential for precise, resource-efficient HPO in classification tasks, complementing the broader applicability of AgentHPO and advancing the paradigm of intelligent, collaborative multi-agent AutoML.
@@ III Methodology
OptiMindTune is conceptualized as a distributed, intelligent multi-agent framework for hyperparameter optimization. The core principle is to decompose the complex HPO problem into manageable sub-tasks, each handled by a specialized agent powered by LLMs, fostering a collaborative ecosystem that collectively drives towards optimal hyperparameter configurations for scikit-learn classifiers.
@@ III-A OptiMindTune Architecture
The OptiMindTune framework comprises three distinct types of AI agents, each with a specific role and responsibilities, interacting through a shared communication mechanism.
  • Recommender Agent: This agent acts as the intelligent explorer and knowledge synthesizer. It analyzes dataset characteristics and reviews past performance history to suggest suitable machine learning models (e.g., from scikit-learn) and their corresponding hyperparameter configurations.
@@
    • Powered by: Large language models for intelligent reasoning, pattern recognition, and suggestion generation based on contextual information.
@@
    • Powered by: Large language models for strategic reasoning, goal-oriented decision-making, and assessing trade-offs.
@@ III-B1 Communication Flow (Optimization Loop)
    • Recommender Suggests: The Recommender Agent analyzes the dataset characteristics and the current optimization history. It then suggests a set of model and hyperparameter configurations, along with its reasoning, to the Decision Agent.
@@
    • Decision Agent Guides: The Decision Agent evaluates the performance metrics reported by the Evaluator Agent and, based on predefined optimization goals and internal reasoning, determines the next course of action. It decides whether to accept the current configuration, continue exploration, or terminate the search process.
@@ IV-A Datasets
We used three standard UCI classification datasets:
  • Breast Cancer Wisconsin (Diagnostic): Binary classification.
  • Iris: Multi-class classification with three classes.
  • Wine: Multi-class classification with three classes.
@@ IV-B Machine Learning Models
  • LogisticRegression: A linear model.
  • RandomForestClassifier: An ensemble tree-based model.
@@ IV-C Baseline Methods
We compared OptiMindTune against:
@@ IV-D Evaluation Metrics
Performance was assessed using:
  • Achieved Objective Value: Mean cross-validated accuracy.
@@ IV-E Experimental Setup
  • Hardware: Consistent hardware was used.
  • Reproducibility: Fixed random seeds ensured consistency.
  • LLM Usage: Gemini 2.0 Flash was used to support agent reasoning.
@@ V-D Discussion of Results
VI Discussion and Future Work
VI-A Reflection on Findings
The experimental results are expected to validate OptiMindTune’s efficacy as a multi-agent framework for intelligent hyperparameter optimization. We anticipate observing that OptiMindTune achieves competitive or superior performance compared to state-of-the-art baselines in terms of objective value, convergence speed, and computational efficiency across a diverse set of datasets and machine learning models. The strength of OptiMindTune lies in its ability to leverage the collective intelligence of specialized agents, each powered by LLMs, allowing for dynamic adaptation to the HPO landscape. The clear division of labor among the Recommender, Evaluator, and Decision Agents, with their sophisticated communication and internal reasoning, facilitates a more comprehensive and adaptive search strategy than monolithic approaches. The explicit real-time logging and conversation tracking also provide unprecedented transparency into the AutoML process.
VI-B Limitations
@@
  • Interpretability of LLM Decisions: While the agents provide reasoning, the internal workings of the LLMs themselves remain a black box. Understanding precisely why an agent makes a certain recommendation or decision can be challenging, impacting the overall interpretability of the HPO process.
@@
VI-C Potential Extensions and Future Work
@@
VII Conclusion
Hyperparameter optimization remains a formidable challenge in machine learning, demanding intelligent and efficient solutions. This paper introduced OptiMindTune, a novel multi-agent framework that reimagines HPO as a collaborative effort among specialized intelligent agents, each powered by large language models. By decoupling responsibilities into distinct roles for recommendation, evaluation, and strategic decision-making, and fostering dynamic interactions, OptiMindTune aims to overcome the limitations of traditional monolithic approaches. The proposed architecture promises enhanced adaptability, robustness, and efficiency in navigating complex hyperparameter landscapes for scikit-learn classifiers. While initial conceptualization and experimental design lay a strong foundation, the future work outlined points towards exciting avenues for extending OptiMindTune to tackle even more challenging optimization problems in the rapidly evolving field of machine learning, especially in the context of deep learning and neural architecture search, leveraging the ever-increasing capabilities of advanced AI models. OptiMindTune represents a significant step towards truly intelligent and autonomous hyperparameter optimization.